### v1

# Improving Parallel Program Performance Through DSL-Driven Code Generation with LLM Optimizers

Abstract
Mapping computations to processors and assigning data to memory are critical for maximizing performance in parallel programming. These mapping decisions are managed through the development of specialized low-level system code, called mappers, crafted by performance engineers. Each mapper is tailored to a specific application and optimized for the underlying machine architecture, a process that requires days of refinement and tuning from an expert. Despite advances in system research, automating mapper generation remains a challenge due to the complexity of making millions of decisions to find the optimal solution and generate the solution as code. We introduce an approach that leverages recent advances in LLM-based optimizers for mapper design. In under ten minutes, our method automatically discovers mappers that surpass human expert designs in scientific applications by up to 1.34×1.34$\times$1.34 × speedup. For parallel matrix multiplication algorithms, our mapper achieves up to 1.31×1.31$\times$1.31 × of the expert-designed solution. To achieve this, we simplify the complexity of low-level code generation by introducing a domain-specific language (DSL) that abstracts the low-level system programming details and defines a structured search space for LLMs to explore. To maximize the application performance, we use an LLM optimizer to improve an agentic system that generates the mapper code. As a result, this approach significantly reduces the workload for performance engineers while achieving substantial performance gains across diverse applications. Finally, our results demonstrate the effectiveness of LLM-based optimization in system design and suggest its potential for addressing other complex system challenges.
1 Introduction
Task-based programming (Slaughter et al., 2015; Bauer et al., 2012; Augonnet et al., 2009; Chamberlain et al., 2007; Kaiser et al., 2014; Heller et al., 2017; Chandra et al., 2001; Duran et al., 2011; Moritz et al., 2018; Barham et al., 2022) has recently emerged as a prominent paradigm in parallel programming. The core idea is to decompose computations into self-contained functions called tasks that do not communicate with other tasks except through their arguments (typically tensors or multi-dimensional arrays). A performance-critical aspect of executing task-based applications is mapping, which involves assigning tasks to processors, mapping data to physical memories, and managing other low-level hardware resources. A concrete mapping policy, implemented as code, is referred to as a mapper. Mappers focus solely on improving performance without altering the correctness of the application’s output. The performance impact of a good versus a poor mapper can be substantial, often resulting in differences of an order of magnitude (Abdel-Gawad et al., 2014; Galvez et al., 2017).
Currently, writing mappers is a manual, labor-intensive, and time-consuming process that requires deep knowledge of the application, the machine, and the low-level system’s C++ mapping APIs. Even for experienced engineers, developing a high-performance mapper can take several days, requiring multiple rounds of refinement (SFX Teixeira et al., 2023).
In this paper, we focus on using LLMs to automatically generate mappers. There are two main challenges in using LLMs for mapper code generation. First, given a specific mapping policy described in natural language, LLMs struggle to generate the corresponding mapper code, which consists of a few hundred lines of low-level C++ code. This code generation task is significantly different from function-level code generation, which is more isolated and modular, where LLMs have demonstrated strong performance (Chen et al., 2021; Austin et al., 2021). In our case, generating C++ mappers requires a deep understanding of low-level systems and broader contextual dependencies in large-scale system code, which is beyond the capabilities of current LLMs (Jimenez et al., 2023; Du et al., 2023; Wang et al., 2024). The second challenge lies in identifying the optimal mapping strategy to maximize application performance. Many design problems in systems are naturally constrained optimization problems. Generating mappers that are syntactically correct is insufficient.
To address the first challenge, we design a Domain-Specific Language (DSL) for mapper generation. The DSL provides a high-level interface that encapsulates all the critical decisions a mapper has to make without interacting with the intricate low-level C++ APIs. The DSL defines a search space in its abstractions, which breaks a monolithic code generation down to generating multiple independent, modular components, which is much easier for code generation due to less contextual dependence between components.
For the second challenge, oirst, structuring the search space through the DSL’s decomposition and modularization of the mapper, and second, applying reinforcement learnformulating the search problem as an optimization problem where the objective is to maximize the performance of the parallel program. Similar approaches have achieved significant results in discovering faster algorithms for matrix multiplication (AlphaTensor) (Fawzi et al., 2022) and generating mathematical proofs (AlphaGeometry) (Trinh et al., 2024), both require searching through a discrete space for an optimal solution.
Recently, a great deal of work has demonstrated that instruction-tuned LLMs have an inherent ability to “optimize” for an objective, such as minimizing a function (Nie et al., 2024; Cheng et al., 2024), prompt optimization (Yang et al., ; Khattab et al., 2023), and molecule design optimization (Yuksekgonul et al., 2024). We propose a fully automated agent-based system to generate mappers and optimize for parallel application performance, illustrated in Figure 1. Our solution allows LLMs to refine their mapping strategies efficiently, reducing the time required to develop high-performance mappers from days to just a few minutes.
Figure 1: We illustrate an agent-based optimization for high-performance mapper code generation, where the mapper is written in a Domain-Specific Language (DSL) and iteratively refined by an agent through performance feedback. Initially, all tasks are mapped to the CPU and system memory. The LLM then improves by mapping computation-intensive tasks to the GPU and GPU’s FrameBuffer memory to minimize task execution time. Finally, tasks that require significant data exchange between the CPU and GPU are mapped to the ZeroCopy memory, enabling shared access and minimizing data movement overhead.
Our experiments show that mappers generated by LLMs not only match but can surpass expert-written mappers, achieving up to 1.34×1.34$\times$1.34 × speedup across 9 benchmarks. Notably, the search process enhances the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×. We also demonstrate, through ablation studies, the critical role of high-quality feedback in guiding LLM optimizers to discover more efficient mapping strategies. Our contributions are as follows:
  1. 1.
    Development of a Domain-Specific Language (DSL) to Simplify Mapper Generation: We design a DSL to address the complexities of direct code generation in low-level systems. This DSL provides a higher-level abstraction, enabling LLMs to generate mapping code more effectively by encapsulating the complexities of low-level C++ APIs.
  2. 2.
    Formulation of Mapper Generation as a Discrete Optimization Problem: We formulate the task of mapper code generation as a discrete optimization problem with the objective of maximizing application performance. Our DSL defines a more structured and constrained search space for LLM to explore.
  3. 3.
    Experimental Validation and Performance Comparison: Our DSL significantly improves the success rate of mapper code generation for a given mapping strategy from 0% in C++ to 80% in our DSL. Leveraging high-quality feedback, our agent-based solution discovers mappers that achieve up to 1.34×1.34$\times$1.34 × speedup compared with expert-designed mappers, enhancing the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×.
2 Related Work
Mapping in Parallel Programming Many parallel programming systems allow users to make their own mapping decisions, such as Legion (Bauer et al., 2012), StarPU (Augonnet et al., 2009; 2010), Chapel (Chamberlain et al., 2007), HPX (Kaiser et al., 2014; Heller et al., 2017), Sequoia (Fatahalian et al., 2006), Ray (Moritz et al., 2018), and Pathways (Barham et al., 2022). Several techniques have been proposed to automate mapping, including machine learning models (O’Boyle et al., 2013; Wang & O’Boyle, 2009), static analysis (Poesia et al., 2017; Ren et al., 2008), and auto-tuning (SFX Teixeira et al., 2023). Unlike previous work, we use an agent-based approach with LLMs, exploring a larger search space for mappers than traditional methods.
Automated Code Generation With the rise of LLMs, various models (Chen et al., 2021; Nijkamp et al., 2022; Li et al., 2023; Ouyang et al., 2022; Wei et al., 2023) have been trained on code, enabling them to assist with programming tasks and automate software development (Shypula et al., 2023; Yang et al., 2024b; Jain et al., 2024). While LLMs have shown success in generating isolated function-level Python code (Chen et al., 2021; Austin et al., 2021), they face challenges when generating system-level C++ code in large repositories due to complex APIs and broader contextual dependencies (Du et al., 2023; Wang et al., 2024). Thus, adopting LLMs to generate code in real-world software remains difficult (Jimenez et al., 2023). Our work addresses this by developing a domain-specific language (DSL) that provides a high-level abstraction to help LLMs generate mapper code more effectively.
Automated System Design Optimization The use of machine learning for optimizing system design has garnered significant attention in recent years. Learned cost models have proven effective in optimizing deep learning operators on GPUs (Zheng et al., 2020a; b; 2022b). Methods in reinforcement learning and multi-armed bandit have been applied to tasks such as chip floorplanning (Mirhoseini et al., 2021) and program autotuning (Ansel et al., 2014), automatic vectorization of programs (Haj-Ali et al., 2020a), and optimizing compiler phase ordering (Haj-Ali et al., 2020b). Previous work primarily focuses on hyperparameter configurations, while our work aims to improve system performance by directly generating mapper code.
Optimization with LLMs
Recently, LLMs have been increasingly applied to optimization problems traditionally found in numerical domains (Boyd & Vandenberghe, 2004). AhmadiTeshnizi et al. (2023) addressed structured optimization challenges, such as mixed-integer linear programming (LP), while Yang et al. (2024a) applied LLMs to unstructured problems, treating them as black-box optimizations. (Nie et al., 2024) highlighted the role of feedback in efficiently finding global minima for black-box functions. Cheng et al. (2024) formally defined optimization problems solvable by LLMs and proposed an optimizer based on execution graphs. More recently, Yuksekgonul et al. (2024) introduced an LLM optimizer inspired by gradient descent. These advancements demonstrate the potential of LLMs to optimize beyond task completion, improving performance metrics across various domains.
3 Mapper Generation Task
Task Definition
The problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012). Mappers determine how tasks are assigned to processors and how data is placed in memory to optimize performance. A well-designed mapper can achieve up to 10$\times$ speedup compared to random mapping strategies. However, mappers are highly context-specific and must be carefully tailored to an application’s input and the machine’s architecture. Finding an optimal mapper is akin to solving a combinatorial optimization problem—a process that is time-consuming, labor-intensive, and traditionally performed by experienced performance engineers. The search space for discovering the best mapping strategy is vast, growing to $2^{14}$ even for the simplest scientific applications (as shown in prior work with a smaller search space than ours (SFX Teixeira et al., 2023)), making it challenging to efficiently explore all possible solutions. Moreover, even with a clear mapping strategy, writing the corresponding mapper requires experts to produce hundreds of lines of low-level C++ code, a task that can take several hours. As a result, the entire performance tuning process can take several days to complete.
Figure 2: Mappers decide the placement of each task in the task graph to processors, the placement of data to memory, and how the iteration space of data is partitioned and mapped to different processors.
Mapping Decisions
Now we elaborate on the decisions that a mapper has to make. The first critical decision is the processor selection for each task, determining whether a task is better suited for GPUs, CPUs, or OpenMP runtime. This choice depends on factors such as task size, GPU memory, and kernel launch latency. For example, tiny tasks that require very little computation may prefer to run on CPUs due to the GPU kernel launch overhead, whereas tasks with large memory footprints might prefer to be assigned to OpenMP or CPU when GPU memory is insufficient.
Next, the memory placement of data across different memory spaces is crucial to performance. A mapper must decide where to place data — in the GPU’s FrameBuffer for faster access, in ZeroCopy memory for shared access between CPU and GPU, or in CPU system memory for large data. Each choice introduces a trade-off between memory access speed, memory usage, and transfer overhead.
The memory layout option determines the optimal memory arrangement for data structures. Depending on the memory access patterns and the underlying hardware, selecting between a Struct of Arrays (SOA) or an Array of Structures (AOS) layout, along with constraints on array ordering (Fortran-order or C-order) and memory alignment, can significantly impact performance due to cache efficiency and data locality. Such choice is task-dependent and processor-dependent.
Finally, mappers need to decide how to perform index mapping. As shown in Figure 2, index mapping controls how the data (or more accurately, parallel for loops of task launches) is partitioned and mapped to the processors. The data (or tensors) can be multi-dimensional, and the distributed machines can also be viewed as a processor space (nodes and processors within the node). Index mapping controls the mapping between the two index spaces. Prior work (Unger et al., 2022; Zheng et al., 2022a) shows that searching over how data is partitioned and mapped can significantly change inter-processor communication volume, which affects performance. Commonly-used index mapping functions are shown in Section A.3.
4 A New Approach: Generation and Optimization in a DSL
4.1 Design a Domain-Specific Language
We designed a domain-specific language for mapper generation. We distill the key performance-critical aspects of mapping into language constructs that are easily expressed in the DSL. By providing a higher-level abstraction than C++ APIs, the DSL simplifies interfacing with LLMs. Also, such a modular design enables LLMs to efficiently address the complex optimization challenges of generating high-performance mappers. To implement our DSL, we develop a compiler that can translate the mapper written in DSL into low-level C++ mapping APIs.
Figure 3(a) presents an example DSL mapper that illustrates the key features of our domain-specific mapping language. By contrast, Figure 3(b) shows the (already simplified) code snippet extracted from a C++ mapper. Notably, the C++ snippet covers only part of the functionality that the IndexMap statement (which we discuss in detail below) in the DSL provides.
While a typical C++ mapper requires approximately 400 lines of code, the corresponding DSL mapper can achieve the same functionality in just about 30 lines on average — reflecting a 14×14$\times$14 × reduction, as detailed in Table 1 (based on the 9 applications used in Section 5.2 and Section 5.3). This substantial reduction in code complexity makes the DSL a more accessible target for LLM-based code generation, as it abstracts away many low-level implementation details.
1# Map task0 to GPU.
2Task task0 GPU;
3
4# Place certain data onto GPU ZeroCopy.
5Region * ghost_region GPU ZCMEM
6
7# Specify layout in memory
8# (aligned to 64 bytes)
9Layout * * * C_order SOA Align==64
10
11# Define a cyclic mapping strategy
12def cyclic(Task task):
13 ip = task.ipoint;
14 mgpu = Machine(GPU);
15 node_idx = ip[0] % mgpu.size[0];
16 gpu_idx = ip[0] % mgpu.size[1];
17 return mgpu[node_idx, gpu_idx];
18
19IndexTaskMap task4 cyclic
(a) An example mapper in DSL
1void slice_task(const Task& task,
2 const SliceTaskInput &input,
3 SliceTaskOutput &output) {
4 vector<Processor> targets =
5 this->select_targets_for_task(ctx, task);
6 DomainT<2> space = input.domain;
7 Point<2> num_points =
8 space.bounds.hi - space.bounds.lo + ones;
9 Rect<2> blocks(zeroes, num_blocks - ones);
10 ...// **126 lines of C++ code ommitted**
11 for (PointInRectIterator<2> it(blocks); it() != NULL; it++) {
12 DomainT<2,coord_t> slice_space;
13 TaskSlice slice;
14 slice.domain = {slice_lo, slice_hi};
15 slice.proc = targets[index++ % targets.size()];
16 output.slices.push_back(slice);
17 }
18}
(b) Simplified code snippet from a C++ mapper
Figure 3: A DSL mapper and a short code snippet from C++ mapper. The C++ mapper has already been simplified, and the shown C++ snippet only contributes to part of the functionality achieved by the IndexTaskMap statement on Line 19 in the DSL mapper.
Next, we explain the DSL’s design, highlighting the immense search space it opens for optimization, which makes generating high-performance mappers a challenging task. The grammar of the DSL is shown in Section A.1.
The Task statement (Line 2) performs processor selection for tasks. Although it may seem straightforward to always prefer GPUs, this decision is complicated by other factors such as limited GPU memory and kernel execution time as explained in Section 3. This is a per-task decision.
The Region statement (Line 5) performs memory placement for data. Possible choices include GPU’s FrameBuffer or ZeroCopy memory, CPU’s System memory or RDMA memory. Such decisions need to be made for each argument of each task, which opens up a huge trade-off space between memory usage, task execution time, and data transfer costs as explained in Section 3.
The Layout statement (Line 9) defines memory layouts, supporting both SOA (Struct of Arrays) and AOS (Array of Structures), as well as array ordering constraints like C_order, F_order, and memory alignment. Optimizing memory layout is crucial for performance, as access patterns vary with tasks, processors, and data structures. This is a per-task, per-data, per-processor decision.
The IndexTaskMap statement (Line 19) enables the mapping of tasks to processors using custom functions, such as cyclic or block mapping strategies. This creates a mapping between the index space defined in application code (e.g., for loops) and the processor space of distributed machines. Many parallel operators (e.g., matrix multiplications) are based on such loops, and deciding how to map them to distributed machines is inherently difficult. The DSL enables users to express arbitrary arithmetic mappings between two index spaces, significantly increasing the complexity of the search space. Additionally, we introduce primitives that transform the processor space, allowing for more flexible and precise index mapping decisions, as detailed in Section A.2.
Application 1 2 3 4 5 6 7 8 9 Avg.
LoC in C++ 347 306 379 447 437 430 428 433 448 406
LoC in DSL 16 14 16 38 38 38 33 38 32 29
LoC Reduction $\times2$ $\times2$ 24×24$\times$24 × 12×12$\times$12 × 12×12$\times$12 × $\times1$ 13×13$\times$13 × $\times1$ 14×14$\times$14 × 14×14$\times$14 ×
Table 1: Lines of Code (LoC) of mapper written in DSL, and LoC reduction compared with C++
4.2 Learning to Generate via Interactive Feedback
Figure 4: We show the high-level design of a mapper agent. The agent takes two inputs: server-related specifications and application-related information. When a mapper is generated by the agent, we run the mapper code on the server and get performance feedback. Then, we use an LLM to update each code block to improve the performance.
1class MapperAgent(trace.Module):
2
3 @trace.bundle(trainable=True)
4 def gen_task_stmt(self, app) -> str:
5 """
6 Generate the policy for placing tasks.
7 Example generated code:
8 Task * GPU; Task task1 CPU;
9 """
10 code = ""
11 for task in app.tasks:
12 proc = random.choice(["GPU","CPU"])
[…]
14 return code
15
16 def forward(self, app):
17 task_stmt = self.gen_task_stmt(app)
18 region_stmt = self.gen_region_stmt(app)
19 layout_stmt = self.gen_layout_stmt(app)
20 ...
21 return task_stmt + region_stmt + ...
(a) The decision procedures in generating a mapper.
1policy = MapperAgent()
2params = policy.parameters()
3optimizer = trace.Optimizer(params)
4
5app = GetApplicationInfo()
6test = GetMapperEvaluator(app)
7
8for i in range(iterations):
9 # Forward pass
10 try:
11 mapper = policy(app)
12 # feedback (str) contains performance
13 feedback = test(mapper)
14 except TraceExecutionError as e:
15 feedback = str(e)
16 target = e.exception_node
17
18 # Backward pass and update
19 optimizer.zero_feedback()
20 optimizer.backward(target, feedback)
21 optimizer.step()
(b) Learning policy using LLM-based optimizers.
Figure 5: Example agent built with the interface offered by Trace. Figure 5(a) shows the mapper agent (each decision is defined as a function). Figure 5(b) shows how we incorporate the feedback from the execution to update the agent. The full agent code is shown in Figure A6.
We conceptualize the mapper generation problem as an online optimization problem. Given a triplet $(\Theta,\omega,{\mathcal{T}})$, where $\Theta$ is the overall set of possible code that can be produced to generate a mapper; $\omega$ defines the optimization objective, which is to create a mapper that maximizes the throughput; and ${\mathcal{T}}:\Theta\rightarrow{\mathcal{F}},{\mathcal{G}}$ is the task that returns $f,g={\mathcal{T}}(\theta)$, where $f$ is the feedback from executing the generated mapper (the performance metric measured after using the generated mapper to serve the application code), and $g$ is the process graph of how the mapper is generated. In our particular setup, the performance of the mapper (throughput) is deterministic because system researchers have carefully controlled all possible randomness in the environment. If the parameter space is numerical, this online optimization problem can be solved through bandit algorithms (Lattimore & Szepesvári, 2020), reinforcement learning (Sutton & Barto, 2018), or Bayesian optimization (Snoek et al., 2012), but these methods are less effective when the parameter is text (mapper code).
In this online optimization problem, we leverage the DSL to bring structure to the parameter space and make the online optimization more efficient. $\theta$ is the program code, $\omega$, and $f$ are expressed as text. We leverage an LLM as an optimizer to solve this optimization problem. LLM has been shown to have some ability to “optimize” by iteratively proposing better solutions when an explicit objective is described in the text. This phenomenon has been observed and leveraged by Yang et al. (2024a); Cheng et al. (2024); Yuksekgonul et al. (2024); Patel et al. (2024) to solve various optimization problems. We design our mapper agent to make various mapping decisions, from selecting processors for tasks to mapping index tasks onto the processor space. A high-level schema of the mapper agent is shown in Figure 4.
Agent Design
We express our mapper agent as a Python program (MapperAgent in Figure 5(a)). This program serves as the initial starting point of the optimization. In this mapper agent, we manually break down the task of generating a monolithic mapper into several independent code segments, allowing the agent to decide what code to generate for each segment. This strategy works because the design of our DSL removes unnecessary dependence between code segments. Our modularization approach mimics the least-to-most prompting (Zhou et al., 2022), and our efforts to use programs as prompts echo the program-as-thought approach (Chen et al., 2022).
Optimization
We illustrate our workflow design in Figure 5, with the full template in Figure A6. Using the framework Trace (Cheng et al., 2024), we only ask the LLM to generate the method decorated with @trace.bundle(trainable=True) at each iteration of optimization. During an optimization step (shown in Figure 5(b)), the LLM receives complete information about how the agent generated the previous mapper, as well as feedback from the system execution. The LLM optimizer needs feedback from the system to perform the optimization. The system feedback can be categorized into three types: (1) Compile Error, indicating a failure of the LLM to generate code capable of producing a valid DSL; (2) Execution Error, such as an application running out of GPU memory; and (3) Performance Metric, which refers to the application’s running time or throughput. The goal of the optimization process is to minimize running time, thereby increasing throughput. Examples are provided in the “System Feedback” column in Table 2.
Mapper System Feedback Enhanced Feedback
Explain Suggest
mapper1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
mapper2 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
mapper3 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
Table 2: System feedback and enhanced feedback for different mappers.
Enhanced Feedback
To accelerate the LLM’s search for high-performance mappers, we introduce enhanced feedback in addition to system feedback. Table 2 provides examples with enhanced feedback, with further examples available in Section A.7. There are two types of enhanced feedback: the first is to explain execution errors, providing more informative failure messages to guide LLMs in generating correct mapper; the second is to suggest modifications to mapping decisions to avoid errors or improve performance. Enhanced feedback is implemented via keyword matching, where system feedback triggers the corresponding explanations and suggestions for mapper improvements. We conduct an ablation study in Section 5.4 to demonstrate the effectiveness of enhanced feedback in our experiments.
5 Evaluation
Experiments are conducted on a GPU cluster where each node has two Intel 10-core E5-2640 v4 CPUs, 256G main memory, and four NVIDIA Tesla P100 GPUs. Regarding the LLM, we use gpt-4o-2024-08-06.
5.1 Effectiveness of the DSL
To evaluate the necessity and effectiveness of our DSL in mapper code generation, we compared the generation of mappers in C++ with DSL. We devised 10 mapping strategies, each described in natural language, to serve as test cases for code generation [^1] An example strategy is “aligning all data to 64 bytes in memory and utilizing Fortran ordering for multi-dimensional data.”. A complete list of these strategies is provided in Section A.9. It is important to note that the objective is not to optimize the performance of any specific application, but rather to assess whether the LLM can accurately generate a C++/DSL mapper based on a given strategy.
We provided the same types of materials in the prompt to ensure a fair comparison. We provided documentation [^2] We used documentation covering all features of DSL, and pre-existing documentation for C++ mapper., example mapper programs, and starting code for both DSL and C++ interfaces. We measured the success rate of generating mappers that can pass compilation and pass test cases for each strategy. For C++ code generation, we enhanced the process by incorporating compiler feedback for additional attempts. The LLM received compiler error messages and iteratively refined the code up to 10 times. We use DSPy (Khattab et al., 2023) to build our generation pipeline.
Code Generation Target Mapping Strategy Success Rate
1 2 3 4 5 6 7 8 9 10
C++ (single trial) 0%
C++ (iterative refine) 0%
DSL (single trial) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 80%
Table 3: Compilation and Strategy Test. We show the success rate for code generation given 10 mapping strategies. Generating DSL code significantly outperforms generating C++ (either without or with compiler feedback). – fails to compile, ✗ compiles but fails the test, and $\checkmark$ passes the test.
ResultTable 3 presents the performance of the LLM in mapper code generation in three settings: C++ single trial (without compiler feedback), C++ with compiler feedback, and DSL. We evaluate the LLM with 10 different mapping strategies. As observed, the DSL approach consistently outperforms the other settings, followed by C++ with compiler feedback, while C++ without feedback performs the worst. These results confirm that the LLM is poor at generating system-level C++ code and highlight the effectiveness and necessity of our DSL design in significantly enhancing the code generation capabilities for mapping.
Failure Case Analysis The compilation errors in C++ mapper generation arise because LLMs are unable to manage the framework-specific contextual dependencies inherent to low-level system software programming. For instance, LLMs generate variable names or references that do not exist within the provided codebase. These issues are particularly challenging to resolve since we cannot supply the full header files to the LLMs due to token limits.
In the cases where the code compiles but fails the test cases, the failures are exclusively in C++. The root cause lies in the complexity of implementing a C++ mapper, which requires a deep understanding and reasoning about the low-level mapping APIs. As outlined in the documentation, multiple APIs must work in concert to accomplish specific tasks, yet LLMs are unable to grasp this level of coordination. While compiler feedback can guide LLMs in avoiding trivial errors (e.g., such as generating non-existent variable names), it cannot bridge the gap in understanding the intricacies of low-level C++ mapping APIs. This limitation explains why all attempts at C++ code generation ultimately fail.
In contrast, only two test cases fail in the DSL context on the single trial, both due to compilation errors stemming from incorrect usage of the DSL’s syntax. This is understandable given that the DSL is a newly designed language with limited training data available for the LLM. Thus, for performance optimization experiments, we do not attempt to have LLMs optimize over C++ mappers, as results sufficiently demonstrate that LLMs struggle to generate C++ code that meets the specifications of what a desired mapper should do, let alone explore a large search space of different mapping strategies.
Figure 6: Normalized throughput for scientific applications comparing expert-written mappers, random mappers, best mappers found by Trace, and the average optimization trajectories of Trace and OPRO in 10 iterations across 5 runs.
5.2 Accelerating Scientific Applications
We use the following three scientific applications as our benchmarks. The circuit simulation benchmark (Bauer et al., 2012) models the behavior of an electrical circuit by simulating currents and voltages across interconnected nodes and wires. The stencil computation benchmark (Van der Wijngaart & Mattson, 2014) simulates a 2D grid where each point’s value is updated based on its neighbors using a stencil pattern. The Pennant benchmark (Ferenbaugh, 2015) models unstructured mesh, Lagrangian staggered-grid hydrodynamics, used for simulating compressible flow.
In this experiment, we compared the performance of different mappers. All mappers are implemented using our DSL. Our goal is to determine whether LLMs can effectively explore the search space of mappers by generating high-performance DSL mappers.
Expert-written mappers were manually developed and optimized by domain experts as part of the application development process. We re-implemented these expert-written C++ mappers using our DSL to establish a ground truth for comparison with our approach. Validation confirmed that the DSL-based mappers achieve performance equivalent to the original C++ mappers, providing a fair basis for evaluating the performance of our generated mappers.
Randomly generated mappers are produced by our MapperAgent with 10 different random seeds and serve as a baseline for this experiment. We ran 10 randomly generated mappers and reported the average performance.
LLM-generated mappers are implemented using Trace (Cheng et al., 2024), which employs LLMs as optimizers. We tested both the Trace and OPRO (Yang et al., ) search algorithms, running 10 iterations for each application. Due to the stochastic nature of LLM outputs, we repeated the optimization process 5 times and averaged the results. We also report the best mapper found by Trace across the 5 runs.
The performance-critical mapping decisions for these applications are the processor type selection, memory placement of data, and the layout selection for data. Other decisions available to mappers do not change the performance of these applications much. The simplest application above (with the smallest search space) is Stencil, which contains 2 tasks and 12 data arguments. Each task and data argument has two choices of placement, together with 4 additional layout choices for each data argument, forming an optimization space of $2^{38}$.
Results
We use normalized throughput as the performance metric in Figure 6, where higher values indicate better performance. The throughput is normalized w.r.t the expert-written mappers. As shown in Figure 6, random mappers are consistently the least effective across all applications, highlighting the importance of mapping decisions on application performance. When comparing the optimization trajectories of Trace and OPRO, Trace performs similarly to OPRO, and is slightly better than OPRO on Pennant. All the best mappers found by Trace can at least match the performance of expert mappers.
Figure 7: Normalized throughput for matrix-multiplication algorithms. We compare expert-written mappers, random mappers, best mappers found by Trace, and the average optimization trajectories of Trace and OPRO in 10 iterations across 5 runs.
Interestingly, the best mapper identified in the Circuit benchmark outperforms the expert mapper by 1.34×1.34$\times$1.34 ×. Upon manual investigation, we observed that the key difference lies in memory placement: the best mapper allocates two data collections to the GPU FrameBuffer memory, whereas the expert mapper places these collections in GPU ZeroCopy memory. This strategy reduces task execution time, despite a slight increase in inter-GPU communication costs, ultimately leading to improved overall performance. For the Pennant benchmark, while there is a minor difference in data collection placement, the final performance results between the two mappers are nearly equivalent.
For each application, the optimization process completes within 10 minutes, significantly reducing mapper development time from days to minutes. This substantial improvement highlights the efficiency of our LLM-enhanced DSL in quickly generating high-performance mappers, offering clear benefits for both developers and application performance.
5.3 Accelerating Matrix Multiplications
We follow the same experimental setup as described in Section 5.2, with the key difference being the focus on matrix multiplication algorithms in this subsection. Different from the applications in Section 5.2, the important mapping decision here is index mapping, which is the choice of how to launch concurrent tasks and distribute the tiles of matrices across multiple GPUs.
In the DSL mapper, achieving a high-performance index mapping requires specifying a function that maps the iteration space (defined by the parallel for loop) to the processor space of the machine (discussed in Section 4.1 IndexTaskMap). Based on our estimates, each algorithm has approximately $10^{9}$ possible choices for index mapping. We show some of the mapping functions in Section A.5
We target 6 different matrix multiplication algorithms: Cannon’s (Cannon, 1969), SUMMA (Van De Geijn & Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik & Demmel, 2011), and COSMA (Kwasniewski et al., 2019). Each algorithm exhibits different performance characteristics and may be the preferred implementation of matrix multiply depending on the target machine and input size. Our goal is to explore and identify better mappings for these algorithms. We elaborate further on these algorithms in Section A.4.
Results
As shown in Figure 7, the algorithm self-specified expert mappers’ compute throughput is normalized to 1.0. Random mappers only yield 2%-40% of the throughput achieved by expert mappers, highlighting the importance of high-quality mapper design. The best mappers found by Trace consistently outperform the expert mappers, achieving speedups ranging from 1.09×1.09$\times$1.09 × to 1.31×1.31$\times$1.31 ×. For both PUMMA and Solomonik’s, the throughput of the best mapper discovered by Trace is significantly higher than the average optimization trajectory across 5 runs. This variability is due to the randomness of LLMs in our experiments, where even small changes in mappers can lead to notable performance differences. When comparing optimization trajectories of Trace and OPRO, Trace shows significant performance gains in SUMMA, PUMMA, Johnson’s, and COSMA, while achieving similar results on other benchmarks.
Finally, we inspected and compared the mapping decisions of the expert mappers with those of the best mappers found through our search. The performance improvements are entirely attributed to more effective index mapping. Index mapping controls the partitioning and distribution of data, in this case, matrices, across the GPUs. The optimized mapping reduces inter-GPU communication and enhances data locality, leading to improved performance in parallel matrix multiplication algorithms.
Figure 8: Comparison of three types of feedback design. System provides only the system execution information (performance numbers or error messages) to the optimizer. Explain provides additional explanations of execution errors. Suggest provides modification proposals to the mappers. All feedback is automatically generated.
5.4 Ablation Study of Feedback
Within the Trace framework, feedback is provided to the LLM optimizer to inform subsequent improvement on mappers. The quality of the feedback directly affects the success of the optimization process. We evaluate three types of feedback to assess how different feedback influences the performance of LLM-based optimizers during mapper search (described in Table 2): (1) system feedback (compile error, execution error, or performance metrics), (2) error explanations, and (3) suggestions for mapper adjustments.
The first feedback message includes only system feedback (labeled System in Figure 8). Next, we evaluate an enhanced feedback message that includes both system feedback and error explanations (labeled System+Explain). Finally, we provide the full feedback message, including all three types (labeled System+Explain+Suggest), corresponding to the Trace results shown in Figure 6 and Figure 7. We evaluate on three benchmarks: the Circuit simulation benchmark and two matrix multiplication algorithms (COSMA and Cannon’s).
As shown in Figure 8, across all three benchmarks, the full feedback message consistently achieves the highest throughput after 10 iterations, followed by the one without guidance. The system-only feedback performs the worst among the three. While the degree of impact from the feedback types varies across benchmarks, the results clearly demonstrate the critical role of high-quality feedback in guiding LLM optimizers to discover more efficient mappers.
6 Conclusion
In this paper, we addressed the challenges of automating mapper generation in task-based programming through the use of LLMs and a Domain-Specific Language (DSL). By designing a high-level DSL, we effectively simplified the complex task of generating low-level C++ mappers, enabling LLMs to handle mapper generation without deep knowledge of system intricacies. Additionally, we formulated the mapper generation task as a discrete optimization problem, leveraging LLM optimizers to explore a structured and constrained search space defined by the DSL. Our experimental results demonstrate the efficacy of this approach, with LLM-generated mappers achieving up to 1.34×1.34$\times$1.34 × speedup over expert-written mappers across nine benchmarks. For matrix multiplication algorithms, we observed a performance boost of up to 1.31×1.31$\times$1.31 ×. These findings show that our LLM-enhanced DSL significantly reduces development time from days to minutes, benefiting both human developers and the performance of parallel applications.
Reproducibility Statement
We provide comprehensive details regarding the setup of our benchmark, ensuring full reproducibility based on the provided information. We are planning to open-source the framework, as well as the code to reproduce all experiment results. We provide detailed documentation of the evaluation process, along with all the prompts that we use. We use gpt-4o-2024-08-06 with a max token of 16383 and default temperature. All software used in our evaluation (e.g., Legion) is open source. We acknowledge the potential difference in performance measurement across different platforms.
Acknowledgments
We thank Yuhui Zhang, Chenyang Yang, Genghan Zhang, Qizheng Zhang, Mohammad R. Fadiheh, Ed Chen, Mert Yuksekgonul, Jiaxin Ge, Tingyu Zhu, Simon Guo and Anne Ouyang for the feedback. We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
Appendix A Appendix
A.1 DSL Grammar
We show the DSL grammar in Figure A1.
[…]
$\mathsf{IndexTaskMap}\;\mathbf{TaskName}\;\mathbf{var}\mathrel{\mid}$
$\mathsf{SingleTaskMap}\;\mathbf{TaskName}\;\mathbf{var}\mathrel{\mid}$
$\mathtt{FuncDef}\mathrel{\mid}\mathtt{TaskMapModifier}$
[…]
$\mathsf{InstanceLimit}\;\mathbf{TaskName}\;\mathbf{int}\mathrel{\mid}\ldots$
[…]
$\mathtt{Expr}\;\mathsf{Op}\;\mathtt{Expr}\mathrel{\mid}(\mathtt{Expr})\mathrel{\mid}\mathtt{Expr}[\mathtt{Expr}]\mathrel{\mid}*\;\mathtt{Expr}$
$\mathtt{Expr}\;\mathtt{?}\;\mathtt{Expr}:\mathtt{Expr}$
Figure A1: Grammar of DSL. A DSL program is a list of statements, each of which controls one aspect of mapping.
A.2 Processor Space Transformation for Index Mapping
Transformation Semantics
$m^{\prime}=m.\mathtt{split}(i,d)$ $m^{\prime}[a_{0},\ldots,a_{n+1}]\coloneq m[b_{0},\ldots,b_{n-1},b_{n}]$
$b_{t}=\begin{cases}a_{t}\qquad&\quad t<i\\ a_{i}+a_{i+1}\cdot d\qquad&\quad t=i\\ a_{t+1}\qquad&\quad t>i\end{cases}$
$m^{\prime}=m.\mathtt{merge}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1},b_{n}]$
$p<q$ $b_{t}=\begin{cases}a_{t}\qquad&\quad t<p\vee p<t<q\\ a_{p}\%m.size[p]\qquad&\quad t=p\\ a_{p}/m.size[p]\qquad&\quad t=q\\ a_{p-1}\qquad&\quad t>q\\ \end{cases}$
$m^{\prime}=m.\mathtt{swap}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
$b_{t}=\begin{cases}a_{q}\qquad&\quad t=p\\ a_{p}\qquad&\quad t=q\\ a_{t}\qquad&\quad t\neq p\land t\neq q\\ \end{cases}$
$m^{\prime}=m.\mathtt{slice}(i,low,high)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
$0\leq low\leq high<m.size[i]$ $b_{t}=\begin{cases}a_{i}+low\qquad&\quad t=i\\ a_{t}\qquad&\quad t\neq i\\ \end{cases}$
Figure A2: Semantics of processor space transformations expressed as mappings from the indices of the transformed processor space to the indices of the original processor space.
A processor space $m$ can be initialized via m = Machine(GPU); as a 2D tuple where the first dimension represents the node index, and the second dimension represents the index of GPU within the node. To better express mapping between index spaces, we define a set of transformation primitives. A transformation primitive is a function of the processor space $m$ that returns a transformed processor space $m^{\prime}$, where $m$ and $m^{\prime}$ are related through the mapping shown in the right-hand side of Figure A2.
Our transformations are inspired by widely used operations for changing the dimensionality of arrays in libraries such as NumPy, but the application to mapping is quite different. We now explain each transformation in detail.
The split transformation takes two arguments: an integer $i$ that indicates the dimension to be split, and the splitting factor $d$. Suppose $m$ is a processor space of size $(8,8)$, then after executing $m^{\prime}=m.\mathtt{split}(0,2)$, $m^{\prime}$ will be a processor space of size $(2,4,8)$. An important property of split and all DSL transformations is that they are invertible. Thus, mappers can work with the transformed space $m^{\prime}$ but DSL can translate such uses back into the original processor space $m$ to identify which concrete processors to use. In this example, $m^{\prime}[j_{0},j_{1},j_{2}]=m[j_{0}+j_{1}\times 2,j_{2}]$.
The merge transformation takes two dimensions of the original processor to be fused as its input. Suppose $m^{\prime}$ is a processor space of size $(2,4,8)$. After applying $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$, $m^{\prime\prime}$ will be a processor space of size $(8,8)$. The processor indexed by $m^{\prime\prime}[j_{0},j_{1}]$ corresponds to $m^{\prime}[j_{0}\%2,j_{0}/2,j_{1}]$.
Transformations can be chained together. Suppose $m$ is a 2D processor space, and we start with $m^{\prime}=m.\mathtt{split}(0,d)$, followed by $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$. The final processor space $m^{\prime\prime}$ is a 2D processor space. Now we will derive the index transformation from $m^{\prime\prime}$ to $m$ by applying the transformation rules from merge transformation and split transformation: $m^{\prime\prime}[j_{0},j_{1}]=m^{\prime}[j_{0}\%d,j_{0}/d,j_{1}]=m[(j_{0}\%d)+(j_{0}/d)\times d,j_{1}]$. The expression $(j_{0}\%d)+(j_{0}/d)\times d$ can be simplified to $j_{0}$ because the division operator /// between two integers rounds to zero. Therefore, $m^{\prime\prime}[j_{0},j_{1}]=m[j_{0},j_{1}]$, showing that the split and merge transformation primitives are inverses of each other.
The swap transformation primitive takes two parameters indicating the dimensions to be swapped and returns a processor space where the two indices of the chosen dimensions are flipped. The swap transformation is often combined with merge: The merge primitive linearizes two dimensions into one, but there is a choice whether to use row-major or column-major iteration order in the linearization. Users can change the iteration order for the merge by swapping the two dimensions.
The slice transformation primitive takes three parameters, the dimension to slice, and the lower bound and upper bound of the dimension. The index mapping rule for the slice transformation is to add a constant shift in the chosen dimension. The slice transformation primitive can be useful if users want to map the iteration space to only part of the original processor space. If two iteration spaces can be executed concurrently in the program, users can map one iteration space to half of the processors and map the other iteration space to the other half of the processors. In this case, the slice transformation allows users to map an iteration space to the selected portion of the processor space.
A.3 Common Index Mapping Functions
We show the common index mapping functions in Figure A3.
Figure A3: Common transformations and index mapping functions. The shaded subarea of the iteration space will be mapped to the shaded processor in the processor space. The transformation code can transform the original $(2,2)$ processor space into the desired processor space. The result processor space will be used for mapping in the user-defined function.
A.4 Parallel Matrix Multiplication Algorithms
2D Algorithms
Cannon’s (Cannon, 1969) introduced a systolic communication pattern with tiled data partitioning for distributed matrix multiplication. PUMMA (Choi et al., 1994) and SUMMA (Van De Geijn & Watts, 1997) extended this approach by supporting non-square matrices and improving communication efficiency through pipelining. They are called 2D algorithms because they partition the matrices into 2D tiles and then map them onto the processor space.
Non-2D Algorithms
Johnson’s (Agarwal et al., 1995) introduced a 3D algorithm that partitions the input matrices into 3D tiles and uses additional memory per processor to reduce communication compared to 2D algorithms. Solomonik’s (Solomonik & Demmel, 2011) balances between 2D and 3D approaches by using extra memory to further minimize communication. COSMA (Kwasniewski et al., 2019) takes a different approach by optimizing the processor grid and parallelization strategy based on the input size and the machine size.
A.5 Index mapping functions used by Matrix Multiplication Algorithms
We show some index mapping functions used by matrix multiplication algorithms in Figure A4
Figure A4: Example mapping functions used by the mappers of matrix multiplication algorithms.
A.6 Explanation of Index Mapping for Solomonik’s Algorithm
Figure A5: Mapper of the Solomonik’s algorithm on a 2-node machine with 4 GPUs per node. The 2D processor space is transformed via the split transformation primitive into a 6D space (visualized as two 3D spaces).
Figure A5 shows a mapper for the Solomonik’s algorithm on a 2-node machine with 4 GPUs per node. The result distribution for the 3D iteration space is that each node will get half of the whole iteration space by partitioning along the x-axis, and the 4 GPUs per node will perform a 2D block distribution over the y-z plane.
There is a dimension mismatch between the iteration space (3D) and the initial processor space (2D). To conduct the desired mapping required by the algorithm, we first apply the split transformation primitive four times (colored as red in the code). We apply the first (resp. last) two split transformations to make the node dimension (resp. GPU dimension) align with the 3D iteration space. We visualize the result 6D processor space as two 3D spaces. The first 3D space (representing the node dimension) is of size $(2,1,1)$ and the second 3D space (representing the GPU dimension) is of size $(1,2,2)$.
A.7 Examples of Feedback Configurations
We give examples for the system feedback and enhanced feedback in Table A1. The enhanced feedback includes explanations of errors and suggestions for mapper modifications.
Mapper System Feedback Enhanced Feedback
Explain Suggest
mapper1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
mapper2 Compile Error: IndexTaskMap’s function undefined N/A Define the IndexTaskMap function first before using it.
mapper3 Compile Error: mgpu not found N/A Include mgpu = Machine(GPU); in the generated code.
mapper4 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
mapper5 Execution Error: DGEMM parameter number 8 had an illegal value Memory layout is unexpected. Adjust the layout constraint.
mapper6 Execution Error: Slice processor index out of bound IndexTaskMap statements cause error. Ensure that the first index of mgpu ends with % mgpu.size[0], and the second element ends with % mgpu.size[1].
mapper7 Execution Error: Assertion ‘event.exists()’ failed InstanceLimit statements cause error. Avoid generating InstanceLimit statements.
mapper8 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
mapper9 Performance Metric: Achieved throughput = 4877 GFLOPS N/A Try using different IndexTaskMap or SingleTaskMap statements to maximize throughput.
Table A1: System feedback and enhanced feedback (error explanations and adjustment suggestions) for different mappers.
A.8 Trace Agent Code
Trace (Cheng et al., 2024) uses Python decorators like @bundle to annotate Python programs. It allows us to design an LLM code generation agent as if we were writing a Python program ourselves. We first set up an end-to-end runnable Python program that can generate a valid mapper program by randomly making decisions over the search space. We show the high-level structure of our Trace Mapper in Figure A6. At each optimization step, Trace will execute DSLMapperGenerator and collect the corresponding execution flow to build up a graph. Then it will make a call to an LLM to perform an update to any function that is decorated with @bundle(trainable=True). The DSLMapperGenerator is structured in the same way as providing a search space specified by the DSL, where an LLM optimizer can make decisions along the pre-designed axes. We note that this type of design is only enabled by recent developments like Trace and is much more challenging to do using older LLM-based frameworks.
1import opto.trace as trace
2
3class MapperAgent(trace.Module):
4 @trace.bundle(trainable=True)
5 def task_decision(self, tasks):
6 ...
7
8 @trace.bundle(trainable=True)
9 def region_decision(self, regions):
10 ...
11
12 @trace.bundle(trainable=True)
13 def layout_decision(self):
14 ...
15
16 @trace.bundle(trainable=True)
17 def instance_limit_decision(self, tasks):
18 ...
19
20 @trace.bundle(trainable=True)
21 def index_task_map_decision(self, index_tasks):
22
23 @trace.bundle(trainable=True)
24 def single_task_map_decision(self, single_tasks):
25 ...
26
27 def generate_mapper(self):
28 """
29 Generate the final mapper code by combining all code statements.
30 """
31 task_statements = self.task_decision(self.tasks)
32 region_statements = self.region_decision(self.regions)
33 layout_statements = self.layout_decision()
34 instance_limit_statements = self.instance_limit_decision(self.tasks)
35 index_task_map_statements = self.index_task_map_decision(self.index_tasks, self.index_task_specification)
36 single_task_statements = self.single_task_map_decision(self.single_tasks)
37
38 code_statements = (
39 task_statements +
40 region_statements +
41 layout_statements +
42 instance_limit_statements +
43 index_task_map_statements +
44 single_task_statements
45 )
46 # Combine all code statements and function definitions into a single string
47 code_list = code_statements
[…]
49 return mapper_code
Figure A6: High-level structure of the Trace-based agent template, where functions annotated with @bundle(trainable=True) define the search space that the LLM optimizer updates during mapper generation. Note: This agent serves as a shared starting point for ALL tasks. For each task, we produce a mapper from this starting agent and then ask LLMs to “optimize” this agent (by changing functions that are trainable) to produce mappers that are optimal for the particular task.
A.9 Mapping Strategies
Strategy 1: Map the tasks of calculate_new_currents, distribute_charge, update_voltages onto GPUs in this way: linearize the 2D GPU processor space into 1D, then perform 1D block mapping from launch domain to the linearized 1D processor space.
1Task * GPU,CPU; # for any task, run on GPU if supported
2Region * *GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
3Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
4
5Layout * * * SOA C_order;
6
7mcpu = Machine(CPU);
8mgpu = Machine(GPU);
9
10========== Above is fixed ==========
11def linearblock(Task task) {
12 return mgpu[task.ipoint[0] / mgpu.size[1], task.ipoint[0] % mgpu.size[1]];
13}
14
15IndexTaskMap calculate_new_currents,distribute_charge,update_voltages linearblock;
Strategy 2: Place ghost/shared regions (rp_shared and rp_ghost) onto GPU zero-copy memory
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6Layout * * * SOA C_order;
7
8mcpu = Machine(CPU);
9mgpu = Machine(GPU);
10
11========== Above is fixed ==========
12
13Region * rp_shared GPU ZCMEM;
14Region * rp_ghost GPU ZCMEM;
Strategy 3: Use Array Of Struct (AOS) data layout for all data instead of the default SOA
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * AOS;
Strategy 4: Use Fortran ordering of data layout for all data instead of the default C order
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * F_order;
Strategy 5: Align all the regions to 64 bytes while using the Fortran ordering of data
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * Align==64 F_order;
Strategy 6 Place the task calculate_new_currents onto CPU
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7
8mgpu = Machine(GPU);
9
10Layout * * * SOA C_order;
11
12========== Above is fixed ==========
13Task calculate_new_currents CPU;
Strategy 7: Collect all the memory used by task calculate_new_currents
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12CollectMemory calculate_new_currents *;
Strategy 8: Ensure that at most 4 tasks of calculate_new_currents can be run at the same time
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12InstanceLimit calculate_new_currents 4;
Strategy 9: Map the second region argument of task distribute_charge onto GPU’s Zero-Copy memory
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12Region distribute_charge 1 GPU ZCMEM;
Strategy 10: Map the tasks of calculate_new_currents,distribute_charge,update_voltages onto GPUs in a 1D cyclic manner: perform a cyclic distribution over both the node and processor dimensions.
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12def cyclic1d(Task task) {
13 ip = task.ipoint;
14 # cyclic over node, cyclic over gpu
15 return mgpu[ip[0] % mgpu.size[0], ip[0] / mgpu.size[0] % mgpu.size[1]];
16}
17
18IndexTaskMap calculate_new_currents,distribute_charge,update_voltages cyclic1d;
A.10 Generated Mapper Examples
Here we provide examples of generated mappers for a subset of problems. The mappers, written in DSL, are produced by the mapper agent. While the LLM is responsible for creating and refining the mapper agent, the agent itself is implemented in Python, and it generates mappers as DSL programs. For the Circuit Simulation benchmark, the optimized mapper (Figure A8) is more concise than the initial version (Figure A7), with an additional constraint for byte alignment in the data layout. In contrast, for Solomonik’s algorithm, the initial mapper is relatively simple (Figure A9), whereas the final optimized mapper adopts a more complex and detailed index mapping strategy (Figure A10).
1Task * GPU,OMP,CPU;
2Task calculate_new_currents GPU;
3Task update_voltages GPU;
4Region * * GPU FBMEM;
5Region * * * SOCKMEM,SYSMEM;
6Region * all_times GPU FBMEM;
7Region * all_nodes GPU FBMEM;
8Region * all_wires GPU FBMEM;
9Region * ghost_ranges GPU FBMEM;
10Region * rp_all_nodes GPU FBMEM;
11Region * all_private GPU FBMEM;
12Region * all_shared GPU FBMEM;
13Region * rp_shared GPU FBMEM;
14Region * rp_wires GPU FBMEM;
15Region * rp_ghost_ranges GPU FBMEM;
16Layout * * * C_order AOS;
17mgpu = Machine(GPU);
18
19m_2d = Machine(GPU);
20def same_point(Task task) {
21 return m_2d[*task.parent.processor(m_2d)];
22}
Figure A7: For the Circuit task, we show the mapper produced by the mapper agent at iteration 2.
1Task * GPU,OMP,CPU;
2Task calculate_new_currents GPU;
3Task update_voltages GPU;
4Region * * GPU FBMEM;
5Layout * * * C_order AOS Align==128;
6mgpu = Machine(GPU);
7
8m_2d = Machine(GPU);
9def same_point(Task task) {
10 return m_2d[*task.parent.processor(m_2d)];
11}
Figure A8: For the Circuit task, we show the mapper produced by the mapper agent at iteration 10.
1Task * GPU,OMP,CPU;
2Region * * GPU FBMEM;
3Region * * * SOCKMEM,SYSMEM;
4Layout * * * F_order SOA;
5mgpu = Machine(GPU);
6
7def block1d(Task task) {
8 ip = task.ipoint;
9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
10}
11
12IndexTaskMap task_2 block1d;
13
14m_2d = Machine(GPU);
15def same_point(Task task) {
16 return m_2d[*task.parent.processor(m_2d)];
17}
Figure A9: For Solomonik’s algorithm, we show the mapper produced by the mapper agent at iteration 2.
1Task * GPU,OMP,CPU;
2Region * * GPU FBMEM;
3Region * * * SOCKMEM,SYSMEM;
4Layout * * * C_order SOA No_Align;
5mgpu = Machine(GPU);
6
7def block1d(Task task) {
8 ip = task.ipoint;
9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
10}
11
12IndexTaskMap task_1 block1d;
13
14def cyclic1d(Task task) {
15 ip = task.ipoint;
16 linearize = ip[0] * 2 + ip[1];
17 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
18}
19
20IndexTaskMap task_1 cyclic1d;
21
22def cyclic2d(Task task) {
23 ip = task.ipoint;
24 linearize = ip[0] + ip[1] * 2;
25 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
26}
27
28IndexTaskMap task_1 cyclic2d;
29
30def linearize3D(Task task) {
31 ip = task.ipoint;
32 linearize = ip[0] + ip[1] + ip[2];
33 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
34}
35
36IndexTaskMap task_1 linearize3D;
37
38def linearize2D(Task task) {
39 ip = task.ipoint;
40 linearize = ip[0] * 2 + ip[2];
41 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
42}
43
44IndexTaskMap task_1 linearize2D;
45IndexTaskMap task_2 block1d;
46IndexTaskMap task_2 cyclic1d;
47IndexTaskMap task_2 cyclic2d;
48IndexTaskMap task_2 linearize3D;
49IndexTaskMap task_2 linearize2D;
50IndexTaskMap task_3 block1d;
51IndexTaskMap task_3 cyclic1d;
52IndexTaskMap task_3 cyclic2d;
53IndexTaskMap task_3 linearize3D;
54IndexTaskMap task_3 linearize2D;
55IndexTaskMap task_5 block1d;
56IndexTaskMap task_5 cyclic1d;
57IndexTaskMap task_5 cyclic2d;
58IndexTaskMap task_5 linearize3D;
59IndexTaskMap task_5 linearize2D;
60
61m_2d = Machine(GPU);
62def same_point(Task task) {
63 return m_2d[*task.parent.processor(m_2d)];
64}
Figure A10: For Solomonik’s algorithm, we show the mapper produced by the mapper agent at iteration 10.