### v2

# Improving Parallel Program Performance with LLM Optimizers via Agent-System Interface

Abstract
Modern scientific discovery increasingly relies on high-performance computing for complex modeling and simulation. A key challenge in improving parallel program performance is efficiently mapping tasks to processors and data to memory, a process dictated by intricate, low-level system code known as mappers. Developing high-performance mappers demands days of manual tuning, posing a significant barrier for domain scientists without systems expertise. We introduce a framework that automates mapper development with generative optimization, leveraging richer feedback beyond scalar performance metrics. Our approach features the Agent-System Interface, which includes a Domain-Specific Language (DSL) to abstract away low-level complexity of system code and define a structured search space, as well as AutoGuide, a mechanism that interprets raw execution output into actionable feedback. Unlike traditional reinforcement learning methods such as OpenTuner, which rely solely on scalar feedback, our method finds superior mappers in far fewer iterations. With just 10 iterations, it outperforms OpenTuner even after 1000 iterations, achieving 3.8×3.8$\times$3.8 × faster performance. Our approach finds mappers that surpass expert-written mappers by up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks while reducing tuning time from days to minutes.
Machine Learning, ICML
1 Introduction
Modern scientific discovery depends on advanced software tools for modeling and simulation (Stocks et al., 2024; Wang et al., 2024; Ltaief et al., 2024). Computational scientists, including physicists, chemists, and biologists, rely on high-performance computing to tackle complex problems. These scientific computations dominate workloads on the world’s most powerful supercomputers (Exascale Computing, ). However, many domain scientists lack expertise in computer science, and therefore having difficulties in optimizing their programs because of the complexity and scale of the underlying machines. Even for experts, finding and fixing performance problems resulting from program modifications or when porting to a new machine is often time-consuming. Any progress on automating performance tuning is of great benefit in this domain.
Task-based programming (Slaughter et al., 2015; Bauer et al., 2012; Augonnet et al., 2009; Chamberlain et al., 2007; Moritz et al., 2018; Barham et al., 2022) has emerged as a promising approach to high performance computing. The paradigm involves decomposing computations into independent tasks that communicate exclusively through their arguments. A key advantage of task-based systems is that the performance tuning problem is factored out into a separate mapping: an assignment of tasks to processors and data to particular memories. High-quality mapping, achieved through a well-designed mapper (implemented as code), can significantly improve performance, often by an order of magnitude (Galvez et al., 2017).
However, currently writing mappers remains a labor-intensive process, as it requires deep knowledge of applications, hardware, and low-level system APIs. In addition, this process is highly application-specific, input-specific, and machine-specific, often taking experts several days of meticulous tuning to achieve high performance. This challenge is especially pronounced for domain scientists, who typically lack the necessary expertise in computer systems and code optimization. Automating mapper development would enable scientists to focus on their own domain of expertise while fully utilizing the capabilities of high-performance computing systems.
Figure 1: Iterative mapper refinement with agent-based generative optimization. The system leverages the Agent-System Interface, which consists of the Domain-Specific Language (DSL) and AutoGuide. The DSL abstracts away the low-level system code, defining a search space for mapping strategies, while AutoGuide interprets execution results into actionable guidance. As iterations progress, the mapper evolves to improve performance.
In this paper, we introduce a system powered by large language models (LLMs) to automate both the generation and optimization of mapper code. The first challenge stems from the complexity of generating mapper code due to the original low-level programming system, which exposes the agent to intricate system APIs, coupled with the problem that raw feedback messages from the system are often uninformative to the agent. The second challenge involves optimizing mapper performance. Specifically, it consists of (1) defining an appropriate search space and (2) devising efficient methods to find optimal mappers, thereby maximizing parallel program performance.
To address the first challenge, we propose an Agent-System Interface (ASI), as shown in Figure 1, an abstraction layer between the agent and the system that simplifies code generation and provides more meaningful feedback to the agent. At the core of ASI is a Domain-Specific Language (DSL), a high-level interface that encapsulates all performance-critical decisions required to generate a mapper. The DSL abstracts away the complexity of low-level system code with a compiler. Additionally, the DSL defines a structured search space, enabling systematic exploration of mapping strategies. We also design and implement the AutoGuide mechanism to interpret raw execution output into informative and actionable guidance. This mechanism allows the agent to iteratively optimize the mapper by leveraging enriched feedback to update its strategy.
For the second challenge, we adopt the generative optimization approach, a recent advance in optimization techniques. Unlike traditional methods such as reinforcement learning (Ansel et al., 2014), which rely solely on scalar rewards, generative optimization can utilize richer forms of feedback, such as error explanations and actionable suggestions expressed in natural language. This agentic optimization workflow has previously proven to be effective across various domains (Nie et al., 2024; Cheng et al., 2024; Yang et al., ; Khattab et al., 2023; Yuksekgonul et al., 2024). Our work is the first to apply such technique to the domain of system optimization.
Our experiments demonstrate that mappers optimized by LLM-powered agents not only match but often surpass expert-written mappers, achieving up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks. Since expert-written mappers set the highest standard, surpassing them is a notable accomplishment. At the same time, our method significantly reduces mapper tuning time from days to minutes, making high-performance mapping more accessible to domain scientists. To further highlight the advantage of generative optimization, we compare it against OpenTuner, a reinforcement learning-based autotuning framework. Our generative optimizer finds mappers $\times1$ faster than OpenTuner when both run for 10 iterations and still maintains a 3.8×3.8$\times$3.8 × advantage even when OpenTuner runs for 1000 iterations. Furthermore, ablation studies underscore the necessity of the agent-system interface design in achieving these performance gains. Our contributions are as follows:
  1. 1.
    Design of an Agent-System Interface: We introduce an abstraction layer that simplifies mapper code generation and provides guidance to the agent. The Domain-Specific Language (DSL) defines a search space, allowing the agent to explore mapping strategies without dealing with low-level system code. AutoGuide interprets raw execution output into targeted feedback, enabling the agent to refine mapper code more effectively.
  2. 2.
    Generative Optimization for Systems: We introduce generative optimization to improve system performance, leveraging richer feedback such as error messages and actionable suggestions in natural language. Unlike reinforcement learning methods like OpenTuner, which rely solely on scalar feedback, our method identifies better mappers in far fewer iterations. With only 10 iterations, it outperforms OpenTuner by 3.8×3.8$\times$3.8 × even after 1000 iterations.
  3. 3.
    Empirical Evaluation of Performance: Our agent-based solution achieves up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks, surpassing expert-written mappers while reducing tuning time from days to minutes. We highlight the critical role of the agent-system interface through ablation studies, demonstrating its impact on achieving the performance gains.
2 Related Work
Mapping in Parallel Programming Many parallel programming systems allow users to make their own mapping decisions, such as Legion (Bauer et al., 2012), StarPU (Augonnet et al., 2009, 2010), Chapel (Chamberlain et al., 2007), HPX (Kaiser et al., 2014; Heller et al., 2017), Sequoia (Fatahalian et al., 2006), Ray (Moritz et al., 2018), TaskFlow (Huang et al., 2021), and Pathways (Barham et al., 2022). Several techniques have been proposed to automate mapping, including machine learning models (O’Boyle et al., 2013; Wang & O’Boyle, 2009), static analysis (Poesia et al., 2017; Ren et al., 2008), reinforcement learning (Ansel et al., 2014; Mirhoseini et al., 2017) and auto-tuning (SFX Teixeira et al., 2023). We use an agent-based approach with LLMs and explore a larger search space for mappers than traditional methods.
Agentic Frameworks Agents powered by Large Language Models (LLMs) play a critical role in decision-making, planning, tool integration, and solving complex problems in dynamic environments (Guo et al., 2024). Many agentic frameworks have been developed (Yao et al., 2022; Wu et al., 2023; Li et al., 2023; Hong et al., 2023), with uses spanning domains such as software engineering (Gur et al., 2023; Yang et al., 2024b; Jin et al., 2024), robotics (Kannan et al., 2024), healthcare (Li et al., 2024), education (Ramirez & Esparrell, 2024), and knowledge engineering (Shao et al., 2024). Our work is the first to apply an agentic workflow to iteratively optimize mapper code, improving the performance of parallel programs.
AI for Systems The application of AI to optimize system design has gained significant traction in recent years. Techniques such as deep learning (Zheng et al., 2020a, b, 2022b) and gradient-boosted trees (Feng et al., 2023) have been used to predict program execution times for performance optimization. Reinforcement learning methods have addressed challenges in chip floorplanning (Mirhoseini et al., 2021), autotuning (Ansel et al., 2014), auto-vectorization (Haj-Ali et al., 2020a), and compiler phase ordering (Haj-Ali et al., 2020b). While previous efforts have predominantly relied on traditional approaches for cost prediction and optimization, our work uses the recent advances in generative optimization to tackle complex system challenges.
Generative Optimization
Recent work has explored the use of LLMs for optimization problems traditionally tackled with numerical methods, including mixed-integer programming (AhmadiTeshnizi et al., 2024a, b) and numerical optimization (Nie et al., 2024). A key advantage of generative optimization is its ability to iteratively refine solutions using diverse forms of feedback. For example, Cheng et al. (2024) applies generative optimization to robotic manipulation and game playing, while Yuksekgonul et al. (2024) optimizes prompts and molecular designs. While reinforcement learning has been applied to system optimization, the potential of LLM-driven optimization in systems remains unexplored. Our work explores whether generative optimization with richer feedback outperforms traditional methods using scalar rewards in system optimization.
3 Problem Definition
Motivation and Challenges
The concrete problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012). Mappers dictate task scheduling and data placement. A well-designed mapper can achieve orders-of-magnitude speedup over naive strategies.
However, automating mapper generation is challenging due to two key factors. First, the complexity of low-level system code. Implementing a mapper requires writing hundreds of lines of intricate C++ code, demanding expertise in system internals. Second, the vast search space of mapping strategies. The search space grows exponentially with the number of tasks and arguments.
Search Space and Performance Impact
As illustrated in Figure A1, the search space of mappers involves multiple decisions, each influencing performance. The first key aspect is processor selection, which determines whether a task runs on GPUs, CPUs, or the OpenMP runtime. This choice depends on factors such as task size, GPU memory capacity, and kernel launch overhead. For instance, small tasks may prefer CPUs due to the overhead of launching GPU kernels, while tasks with large memory footprints may run on CPUs when GPU memory is insufficient.
Another crucial dimension is memory placement, which dictates where data is stored. A mapper must decide whether to place data in the GPU’s FrameBuffer for fast access, ZeroCopy memory for CPU-GPU sharing, or CPU system memory for more available storage. Each option presents trade-offs between access speed, memory usage, and data transfer overhead.
Additionally, memory layout further expands the search space, with decisions on Struct of Arrays (SOA) vs. Array of Structures (AOS), data ordering (Fortran-order vs. C-order), and alignment constraints (e.g., 128-byte alignment) significantly affecting cache efficiency and performance.
Finally, an important idiom in high-performance computing is launching tasks over partitioned data. Index mapping determines how data partitions and task executions are distributed across multiple processors. For consistency, we can represent data partitioning as a tensor of data partitions, the machine as a tensor of processors, and tasks operating on the partitioned data as a tensor of tasks. The way data and task indices are mapped to processor indices affects inter-processor communication, a key factor in performance (Unger et al., 2022; Zheng et al., 2022a).
4 Our Approach: Agent-System Interface
1# Map task0 to GPU.
2Task task0 GPU;
3
4# Place certain data onto GPU ZeroCopy.
5Region * ghost_region GPU ZCMEM
6
7# Specify layout in memory
8# (aligned to 64 bytes)
9Layout * * * C_order SOA Align==64
10
11# Define a cyclic mapping strategy
12def cyclic(Task task):
13 ip = task.ipoint;
14 mgpu = Machine(GPU);
15 node_idx = ip[0] % mgpu.size[0];
16 gpu_idx = ip[0] % mgpu.size[1];
17 return mgpu[node_idx, gpu_idx];
18
19IndexTaskMap task4 cyclic
(a) An example mapper in Domain-Specific Language (DSL)
1void slice_task(const Task& task,
2 const SliceTaskInput &input,
3 SliceTaskOutput &output) {
4 vector<Processor> targets =
5 this->select_targets_for_task(ctx, task);
6 DomainT<2> space = input.domain;
7 Point<2> num_points =
8 space.bounds.hi - space.bounds.lo + ones;
9 Rect<2> blocks(zeroes, num_blocks - ones);
10 ... // 126 lines of C++ code omitted here
11 for (PointInRectIterator<2> it(blocks); it() != NULL; it++)
12 {
13 DomainT<2,coord_t> slice_space;
14 TaskSlice slice;
15 slice.domain = {slice_lo, slice_hi};
16 slice.proc = targets[index++ % targets.size()];
17 output.slices.push_back(slice);
18 }
19}
(b) Code snippet from a C++ mapper
Figure 2: Comparison of a DSL mapper and a C++ mapper. The DSL’s declarative, high-level design abstracts away the complexity of low-level C++ code, serving as the core of the Agent-System Interface. The highlighted boxes illustrate how the same functionality, which requires extensive C++ system code, can be expressed concisely in just a few lines in DSL.
4.1 Domain-Specific Language Design
A key challenge in automating mapper generation with a coding agent is the complexity of low-level system code, which requires intricate C++ implementations. To address this, we design a high-level Domain-Specific Language (DSL) as the core of our Agent-System Interface (ASI). The DSL provides a structured search space for mapping strategies while abstracting away low-level implementation details. Unlike C++, which demands imperative specifications of mapping policies, our DSL adopts a declarative design, allowing users to specify what to achieve rather than how to implement it. Most critically, the DSL separates concerns, enabling multiple aspects of mapping decisions to be expressed independently rather than being entangled in low-level system APIs. This design reduces code complexity and naturally provides a search space for the agent to explore. To implement it, we develop a compiler that translates DSL into the low-level C++ APIs.
As illustrated in Figure 2, the complexity of DSL code is significantly lower than that of C++. Figure 2(a) provides an example of a DSL mapper, highlighting the key features of our DSL. In contrast, Figure 2(b) shows a snippet from a C++ mapper, emphasizing the intricacy of low-level implementation details. According to Table A1, using the DSL results in an average lines of code reduction of 14×14$\times$14 ×. This substantial reduction makes DSL a more suitable target for LLM code generation, as it abstracts away the complexities inherent in low-level systems. As we will show in Section 5.2, LLMs generate DSL code more effectively, despite DSL having no examples in LLM training corpora, whereas C++ is widely represented.
Next, we describe the DSL’s design, emphasizing its declarative nature and structured search space. Section 3 details the performance impact of each decision.
The Task statement (Line 2) defines processor selection for each task, choosing between CPU, GPU, or OpenMP. Line 2 specifies that instances of task0 should run on GPUs. This decision is made per task; note that the search space expands exponentially with the number of tasks.
The Region statement (Line 5) controls memory placement for data arguments. Line 5 specifies that all tasks using ghost_region should place the data in GPU ZeroCopy memory. Other choices include GPU FrameBuffer memory and CPU System Memory. This decision is made per task and per argument, causing the search space to grow exponentially.
The Layout statement (Line 9) defines memory layouts. Line 9 enforces a C_order axis ordering, an SOA layout, and a 64-byte memory alignment for all data used by all tasks mapped to all processors. Alternative choices include F_order, AOS, and various alignment strategies. This is a per-task, per-data, per-processor decision.
The IndexTaskMap statement (Line 19) controls index mapping using a customized function. Line 12 defines the mapping function that establishes the correspondence between two index spaces: the task index space (represented by task.ipoint) defined in the application code (e.g., for loops) and the processor space of the distributed machine (represented by Machine(GPU)). The DSL allows users to express arbitrary arithmetic mappings between the two index spaces. This decision applies to each task group launched by parallel for loops.
4.2 Generative Optimization via AutoGuide
Figure 3: Agent Optimization Process. The mapper agent takes server specifications and application-specific information as input, generates mapper code, and executes it alongside the application on the server. Raw execution feedback is enriched using the AutoGuide mechanism and iteratively refined by an LLM optimizer to improve performance.
Case Raw Execution Output AutoGuide
Explain Suggest
Case 1 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
Case 2 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
Table 1: AutoGuide Feedback Mechanism. The AutoGuide mechanism interprets raw execution output from the runtime system, providing more informative error explanations and suggestions for mapper modifications. It is implemented via keyword matching. Additional examples are shown in Table A2.
We formulate mapper generation as an online optimization problem. Given a triplet $(\Theta,\omega,\mathcal{T})$, where $\Theta$ is a set of possible mappers, $\omega$ is an optimization objective, and $\mathcal{T}$ is a function that takes a mapper $\theta\in\Theta$ as input, $(f,g)=\mathcal{T}(\theta)$ and returns $f$, the feedback from executing the mapper (i.e., the measured performance after running the application code with the generated mapper), and $g$, the process graph tracing how the mapper was generated. In our setup, mapper performance is deterministic, as we carefully control all sources of randomness in the environment. If the parameter space were numerical, this online optimization problem could be addressed using bandit algorithms (Lattimore & Szepesvári, 2020), reinforcement learning (Sutton & Barto, 2018), or Bayesian optimization (Snoek et al., 2012), but these methods are less efficient when the parameter search space is large and discrete (i.e., text).
In this online optimization problem, we leverage the DSL to structure the parameter space to improve the efficiency of optimization. Here, $\theta$ represents the program code, while $\omega$ and $f$ are expressed as text. We adopt generative optimization, leveraging LLMs as optimizers given the objective in text form. This emergent optimization behavior has been recently observed and applied across various domains (Yang et al., 2024a; Cheng et al., 2024; Yuksekgonul et al., 2024; Patel et al., 2024).
Optimization Process
We present the optimization process in Figure 3. The agent takes two inputs: server specifications (e.g., CPU/GPU counts) and application information (e.g., task lists, data arguments). It generates mapper code, which is executed alongside the application code on the server. Raw execution feedback from the runtime is augmented with the AutoGuide mechanism and fed back to the LLM, iteratively refining the agent for improved mapper code generation.
Coding Agent
Our mapper agent improves mapping decisions by iteratively generating DSL code. A high-level schema of the mapper agent is shown in Figure 3. The mapper agent is implemented as a Python program in the Trace (Cheng et al., 2024) framework, where we decompose the task of generating a monolithic mapper into independent code segments. This decomposition allows the agent to decide what code to generate for each segment separately. This approach is effective because our DSL design eliminates unnecessary dependencies between mapping decisions. Our modularization strategy aligns with least-to-most prompting (Zhou et al., 2022).
AutoGuide
The AutoGuide feedback mechanism is designed based on three key motivations: (1) generative optimization benefits from natural language feedback rather than relying solely on scalar values, (2) raw execution output from the runtime system is often too uninformative to effectively guide the agent’s decisions, and (3) domain heuristics known to systems researchers can be naturally expressed in language (e.g., most tasks run faster on GPUs than CPUs). To address these needs, AutoGuide helps the agent by explaining opaque error messages and suggesting mapper modifications. As shown in Table 1, it interprets uninformative execution output into actionable insights, with additional examples in Section A.4. The implementation relies on keyword matching over the raw execution output. An ablation study in Section 5.3 demonstrates its effectiveness in our experiments.
5 Evaluation
Experiments are conducted on one node with two Intel 10-core E5-2640 v4 CPUs, 256G main memory, and four NVIDIA Tesla P100 GPUs. We use gpt-4o-2024-08-06.
5.1 Speedup of Application Performance
Figure 4: Performance Comparison. Normalized throughput for 9 benchmarks, comparing expert mappers, random mappers, the average optimization trajectories of Trace, OPRO, and OpenTuner in 10 iterations across 5 runs, and the best mappers found by Trace.
Figure 5: Comparison of Trace (generative optimizer) and OpenTuner (traditional RL) over 1K iterations (averaged across all 9 benchmarks)
We evaluate our approach using 9 benchmarks. Circuit (Bauer et al., 2012) is a simulation benchmark that models electrical circuit behavior by simulating currents and voltages across interconnected nodes and wires. Stencil (Van der Wijngaart & Mattson, 2014) simulates a 2D grid where each point’s value is updated based on a stencil pattern determined by its neighbors. Pennant (Ferenbaugh, 2015) models unstructured mesh Lagrangian staggered-grid hydrodynamics, commonly used for simulating compressible flow. The remaining six benchmarks – Cannon’s (Cannon, 1969), SUMMA (Van De Geijn & Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik & Demmel, 2011), and COSMA (Kwasniewski et al., 2019) – are widely studied parallel matrix multiplication algorithms, which we discuss in more detail in Section A.3.
In this experiment, we evaluate the performance of the mappers with the following baselines.
Expert-Written Mappers. These mappers are manually developed by domain scientists who spend years mastering computational science. Writing mappers in parallel programming frameworks is another challenge, and tuning them for specific applications can take days.
Randomly Generated Mappers. These mappers were randomly generated with 10 different random seeds, sampling from the entire search space of each application. We report the average performance.
Agent-Optimized Mappers. Using Trace (Cheng et al., 2024), we evaluated the Trace and OPRO (Yang et al., ) search algorithms, running 10 iterations per application. To account for stochastic output, we repeated the process 5 times and report the average. The best mapper from Trace across runs is also reported.
OpenTuner Mappers. OpenTuner (Ansel et al., 2014) is a program autotuning framework that uses reinforcement learning to optimize performance based on scalar feedback. We provided execution time as feedback, with a high penalty for failures.
Results
We use normalized throughput as the performance metric in Figure 4, where higher values indicate better performance. The throughput is normalized relative to the expert-written mappers. All the best mappers found by Trace can match or surpass the expert-written mappers, underscoring the effectiveness of agent-based generative optimizer. Random mappers consistently exhibit low performance across all applications, emphasizing the critical role of mapping decisions. When comparing optimization trajectories, Trace performs similarly to OPRO, and significantly outperforms OpenTuner.
To further compare the agent-based optimizer with traditional reinforcement learning, we extended OpenTuner’s optimization iterations from 10 to 1000, as shown in Figure 5, where the x-axis is the log-scale of iterations and the y-axis represents relative throughput (averaged across all 9 benchmarks). Notably, Trace achieves a 3.8×3.8$\times$3.8 × speedup over OpenTuner even when OpenTuner is run for 1000 iterations. When both are limited to 10 iterations, Trace outperforms OpenTuner by $\times1$, demonstrating its ability to quickly identify high-performance mappings. This highlights the superiority of Trace (generative optimizer) over OpenTuner (traditional reinforcement learning). Moreover, Trace completes the entire optimization process in just 10 minutes per application, reducing mapper development time from days to minutes.
Case Analysis
The largest performance gain achieved by Trace over the expert mapper is observed in Circuit, with a speedup of 1.34×1.34$\times$1.34 ×. This improvement is primarily due to memory placement: the best mapper allocates two data collections to GPU FrameBuffer memory, while the expert mapper places them in GPU ZeroCopy memory. Despite a slight increase in inter-GPU communication costs, Trace reduces task execution time due to faster memory access, resulting in higher overall performance. For matrix-multiplication algorithms, the greatest speedup is seen in COSMA, with Trace achieving a 1.31×1.31$\times$1.31 × speedup over the expert mapper. This is attributed to Trace’s more efficient index mapping functions, which reduce inter-GPU communication by better distributing partitioned submatrices across GPUs. For additional context, examples of Trace mappers are presented in Section A.7.
Code Generation Target Mapping Strategy Success Rate
1 2 3 4 5 6 7 8 9 10
C++ (single trial) 0%
DSL (single trial) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 80%
C++ (iterative refine) 0%
DSL (iterative refine) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 100%
Table 2: Code Generation Success Rates. Success rates for generating code across 10 mapping strategies described in natural language. The test evaluates whether the generated code compiles and passes execution tests. Generating DSL code significantly outperforms generating C++ for both settings. Symbols indicate results: – fails to compile, ✗ compiles but fails the test, and $\checkmark$ passes the test.
Figure 6: Comparison of different feedback designs. 0-Shot and 5-Shot are baselines. Execution provides only the raw execution output as feedback. Explain provides additional explanations of execution errors. Suggest offers mapper modification suggestions. All feedback is automatically generated.
5.2 Ablation Study of DSL for Code Generation
In Section 5.1, we demonstrate the overall effectiveness of our approach. Here, we conduct an ablation study on the DSL, the core of the Agent-System Interface. Since successful generation is the foundation of optimization, this subsection focuses on how well the DSL helps LLMs generate correct mappers compared to C++, rather than directly optimizing performance.
Experiment Setup
We designed 10 mapping strategies, described in natural language, to evaluate whether LLMs can generate correct code in both the DSL and the original low-level C++. The strategies are detailed in Section A.6. To ensure a fair comparison, identical prompt materials (documentation, examples, and starting code) were provided for both the DSL and C++. Success rates are measured based on whether the generated code passes predefined test cases, with results reported for single trials and iterative refinement, where the LLM is allowed up to 10 iterations to improve the code using compiler feedback. The evaluation is conducted with the DSPy (Khattab et al., 2023) framework.
Results
Table 2 shows that DSL achieves significantly higher generation success rates than C++ in both the single-trial and iterative refinement settings. This demonstrates the effectiveness of DSL’s design in abstracting system complexity and providing a high-level interface that enables LLMs to tackle complex system challenges in code generation. Incorporating iterative refinement with compiler feedback further improves success rates, resolving three compilation errors in C++ and two in the DSL. However, the gap between DSL mappers and C++ mappers remains substantial. Notably, these results are striking given that the DSL is a low-resource language with no pre-training or fine-tuning data, while C++ code is widely present in LLM training corpora.
Analysis
LLMs perform better with the DSL for two reasons. First, the semantic gap between natural language and code is smaller with the DSL than with C++. For example, writing a mapper to “align all data to 64 bytes in memory and use Fortran ordering” requires one line Layout * * * Align==64 F_order in the DSL because of its declarative design. In contrast, the C++ mapping API requires a sequence of operations to enforce alignment and ordering, which widens the semantic gap. Second, the DSL reduces the amount of code. As shown in Table A1, LLMs achieve an average reduction of 14×14$\times$14 × in lines of code, simplifying code generation. These results underscore the importance of a high-level agent-system interface.
5.3 Ablation Study of the AutoGuide Feedback
The AutoGuide mechanism provides enriched feedback to the agentic optimizer. We compare with alternative feedback designs.
Experiment Setup
We compare the following baselines. 0-shot and 5-shot have no feedback, allowing the LLM to generate once with either 0 or 5 examples provided. Execution only provides raw execution feedback, Explain offers additional explanations for execution errors, and Suggest offers mapper modification suggestions. The Trace trajectory shown in Figure 4 uses the full AutoGuide mode with all Execution+Explain+Suggest. As an ablation study, we evaluate 3 benchmarks.
Results and Analysis
Figure 6 demonstrates that the full feedback mechanism consistently outperforms all reduced feedback variants. The 0-shot and 5-shot results perform the worst, underscoring the importance of feedback-based iterative refinement. This highlights the value of an agentic workflow, showing that performance improvements are not solely driven by prompting the LLM but are a direct result of the iterative refinement inherent in the workflow design.
6 Conclusion
In this paper, we introduced a system that leverages LLMs to automate mapper generation and optimization. The Agent-System Interface (ASI) simplifies code generation with a Domain-Specific Language (DSL), which abstracts away the low-level complexity of system code, and enriches execution feedback through AutoGuide, which interprets raw execution output into actionable guidance. We adopted generative optimization, allowing LLMs to refine mappers using rich textual feedback beyond scalar metrics. Unlike RL-based methods like OpenTuner, which rely on numerical rewards, our approach incorporates error explanations and targeted suggestions, accelerating search efficiency. Experiments show that agent-generated mappers outperform expert-written ones, achieving up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks. Our method, running only 10 iterations, maintains a 3.8×3.8$\times$3.8 × advantage over OpenTuner even after 1000 iterations. By reducing mapper development time from days to minutes, our approach benefits computational scientists and demonstrates the effectiveness of generative optimization in system design.
Acknowledgements
We thank Yuhui Zhang, Chenyang Yang, Genghan Zhang, Qizheng Zhang, Mohammad R. Fadiheh, Ed Chen, Mert Yuksekgonul, Jiaxin Ge, Xinyan Hu, Tingyu Zhu, Simon Guo, Anne Ouyang, and Adhish Ramkumar for the feedback. We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Illustration for Mapping
We show an illustration for mapping in Figure A1.
Figure A1: Mappers decide the placement of each task in the task graph to processors, the placement of data to memory, and how the iteration space of data is partitioned and mapped to different processors.
A.2 Tables for Lines of Code Comparison
We show the lines of code comparison between DSL and C++ mappers in Table A1.
Application 1 2 3 4 5 6 7 8 9 Avg.
LoC in C++ 347 306 379 447 437 430 428 433 448 406
LoC in DSL 16 14 16 38 38 38 33 38 32 29
LoC Reduction $\times2$ $\times2$ 24×24$\times$24 × 12×12$\times$12 × 12×12$\times$12 × $\times1$ 13×13$\times$13 × $\times1$ 14×14$\times$14 × 14×14$\times$14 ×
Table A1: Lines of Code (LoC) comparison between DSL and C++ mappers. The DSL achieves a 14×14$\times$14 × average reduction in LoC, making it a more suitable target for LLM code generation.
A.3 Parallel Matrix Multiplication Algorithms
2D Algorithms
Cannon’s (Cannon, 1969) introduced a systolic communication pattern with tiled data partitioning for distributed matrix multiplication. PUMMA (Choi et al., 1994) and SUMMA (Van De Geijn & Watts, 1997) extended this approach by supporting non-square matrices and improving communication efficiency through pipelining. They are called 2D algorithms because they partition the matrices into 2D tiles and then map them onto the processor space.
Non-2D Algorithms
Johnson’s (Agarwal et al., 1995) introduced a 3D algorithm that partitions the input matrices into 3D tiles and uses additional memory per processor to reduce communication compared to 2D algorithms. Solomonik’s (Solomonik & Demmel, 2011) balances between 2D and 3D approaches by using extra memory to further minimize communication. COSMA (Kwasniewski et al., 2019) takes a different approach by optimizing the processor grid and parallelization strategy based on the input size and the machine size.
A.4 Examples of Feedback Configurations
We give examples for the raw execution output and enriched feedback in Table A2. The enhanced feedback includes explanations of errors and suggestions for mapper modifications.
Case Raw Execution Output AutoGuide
Explain Suggest
case1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
case2 Compile Error: IndexTaskMap’s function undefined N/A Define the IndexTaskMap function first before using it.
case3 Compile Error: mgpu not found N/A Include mgpu = Machine(GPU); in the generated code.
case4 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
case5 Execution Error: DGEMM parameter number 8 had an illegal value Memory layout is unexpected. Adjust the layout constraint.
case6 Execution Error: Slice processor index out of bound IndexTaskMap statements cause error. Ensure that the first index of mgpu ends with % mgpu.size[0], and the second element ends with % mgpu.size[1].
case7 Execution Error: Assertion ‘event.exists()’ failed InstanceLimit statements cause error. Avoid generating InstanceLimit statements.
case8 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
case9 Performance Metric: Achieved throughput = 4877 GFLOPS N/A Try using different IndexTaskMap or SingleTaskMap statements to maximize throughput.
Table A2: Raw execution output and AutoGuide (error explanations and adjustment suggestions) for different cases.
A.5 Trace Agent Code
Trace (Cheng et al., 2024) uses Python decorators like @bundle to annotate Python programs. It allows us to design an LLM code generation agent as if we were writing a Python program ourselves. We first set up an end-to-end runnable Python program that can generate a valid mapper program by randomly making decisions over the search space. We show the high-level structure of our Trace Mapper in Figure A3. Figure A2 shows how we incorporate the feedback from the execution to update the agent. At each optimization step, Trace will execute DSLMapperGenerator and collect the corresponding execution flow to build up a graph. Then it will make a call to an LLM to perform an update to any function that is decorated with @bundle(trainable=True). The DSLMapperGenerator is structured in the same way as providing a search space specified by the DSL, where an LLM optimizer can make decisions along the pre-designed axes. We note that this type of design is only enabled by recent developments like Trace and is much more challenging to do using older LLM-based frameworks.
1policy = MapperAgent()
2params = policy.parameters()
3optimizer = trace.Optimizer(params)
4
5app = GetApplicationInfo()
6test = GetMapperEvaluator(app)
7
8for i in range(iterations):
9 # Forward pass
10 try:
11 mapper = policy(app)
12 # feedback (str) contains performance
13 feedback = test(mapper)
14 except TraceExecutionError as e:
15 feedback = str(e)
16 target = e.exception_node
17
18 # Backward pass and update
19 optimizer.zero_feedback()
20 optimizer.backward(target, feedback)
21 optimizer.step()
Figure A2: We show how we use Trace to incorporate the feedback from the execution to update the agent, with a Pytorch-like syntax.
1import opto.trace as trace
2
3class MapperAgent(trace.Module):
4 @trace.bundle(trainable=True)
5 def task_decision(self, tasks):
6 ...
7
8 @trace.bundle(trainable=True)
9 def region_decision(self, regions):
10 ...
11
12 @trace.bundle(trainable=True)
13 def layout_decision(self):
14 ...
15
16 @trace.bundle(trainable=True)
17 def instance_limit_decision(self, tasks):
18 ...
19
20 @trace.bundle(trainable=True)
21 def index_task_map_decision(self, index_tasks):
22
23 @trace.bundle(trainable=True)
24 def single_task_map_decision(self, single_tasks):
25 ...
26
27 def generate_mapper(self):
28 """
29 Generate the final mapper code by combining all code statements.
30 """
31 task_statements = self.task_decision(self.tasks)
32 region_statements = self.region_decision(self.regions)
33 layout_statements = self.layout_decision()
34 instance_limit_statements = self.instance_limit_decision(self.tasks)
35 index_task_map_statements = self.index_task_map_decision(self.index_tasks, self.index_task_specification)
36 single_task_statements = self.single_task_map_decision(self.single_tasks)
37
38 code_statements = (
39 task_statements +
40 region_statements +
41 layout_statements +
42 instance_limit_statements +
43 index_task_map_statements +
44 single_task_statements
45 )
46 # Combine all code statements and function definitions into a single string
47 code_list = code_statements
[…]
49 return mapper_code
Figure A3: High-level structure of the Trace-based agent template, where functions annotated with @bundle(trainable=True) define the search space that the LLM optimizer updates during mapper generation. Note: This agent serves as a shared starting point for ALL tasks. For each task, we produce a mapper from this starting agent and then ask LLMs to “optimize” this agent (by changing functions that are trainable) to produce mappers that are optimal for the particular task.
A.6 Mapping Strategies
Strategy 1: Map the tasks of calculate_new_currents, distribute_charge, update_voltages onto GPUs in this way: linearize the 2D GPU processor space into 1D, then perform 1D block mapping from launch domain to the linearized 1D processor space.
1Task * GPU,CPU; # for any task, run on GPU if supported
2Region * *GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
3Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
4
5Layout * * * SOA C_order;
6
7mcpu = Machine(CPU);
8mgpu = Machine(GPU);
9
10========== Above is fixed ==========
11def linearblock(Task task) {
12 return mgpu[task.ipoint[0] / mgpu.size[1], task.ipoint[0] % mgpu.size[1]];
13}
14
15IndexTaskMap calculate_new_currents,distribute_charge,update_voltages linearblock;
Strategy 2: Place ghost/shared regions (rp_shared and rp_ghost) onto GPU zero-copy memory
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6Layout * * * SOA C_order;
7
8mcpu = Machine(CPU);
9mgpu = Machine(GPU);
10
11========== Above is fixed ==========
12
13Region * rp_shared GPU ZCMEM;
14Region * rp_ghost GPU ZCMEM;
Strategy 3: Use Array Of Struct (AOS) data layout for all data instead of the default SOA
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * AOS;
Strategy 4: Use Fortran ordering of data layout for all data instead of the default C order
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * F_order;
Strategy 5: Align all the regions to 64 bytes while using the Fortran ordering of data
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * Align==64 F_order;
Strategy 6: Place the task calculate_new_currents onto CPU
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7
8mgpu = Machine(GPU);
9
10Layout * * * SOA C_order;
11
12========== Above is fixed ==========
13Task calculate_new_currents CPU;
Strategy 7: Collect all the memory used by task calculate_new_currents
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12CollectMemory calculate_new_currents *;
Strategy 8: Ensure that at most 4 tasks of calculate_new_currents can be run at the same time
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12InstanceLimit calculate_new_currents 4;
Strategy 9: Map the second region argument of task distribute_charge onto GPU’s Zero-Copy memory
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12Region distribute_charge 1 GPU ZCMEM;
Strategy 10: Map the tasks of calculate_new_currents,distribute_charge,update_voltages onto GPUs in a 1D cyclic manner: perform a cyclic distribution over both the node and processor dimensions.
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12def cyclic1d(Task task) {
13 ip = task.ipoint;
14 # cyclic over node, cyclic over gpu
15 return mgpu[ip[0] % mgpu.size[0], ip[0] / mgpu.size[0] % mgpu.size[1]];
16}
17
18IndexTaskMap calculate_new_currents,distribute_charge,update_voltages cyclic1d;
A.7 Generated Mapper Examples
Here we provide examples of generated mappers for a subset of problems. The mappers, written in DSL, are produced by the mapper agent. While the LLM is responsible for creating and refining the mapper agent, the agent itself is implemented in Python, and it generates mappers as DSL programs. For the Circuit Simulation benchmark, the optimized mapper (Figure A5) is more concise than the initial version (Figure A4), with an additional constraint for byte alignment in the data layout. In contrast, for Solomonik’s algorithm, the initial mapper is relatively simple (Figure A6), whereas the final optimized mapper adopts a more complex and detailed index mapping strategy (Figure A7).
1Task * GPU,OMP,CPU;
2Task calculate_new_currents GPU;
3Task update_voltages GPU;
4Region * * GPU FBMEM;
5Region * * * SOCKMEM,SYSMEM;
6Region * all_times GPU FBMEM;
7Region * all_nodes GPU FBMEM;
8Region * all_wires GPU FBMEM;
9Region * ghost_ranges GPU FBMEM;
10Region * rp_all_nodes GPU FBMEM;
11Region * all_private GPU FBMEM;
12Region * all_shared GPU FBMEM;
13Region * rp_shared GPU FBMEM;
14Region * rp_wires GPU FBMEM;
15Region * rp_ghost_ranges GPU FBMEM;
16Layout * * * C_order AOS;
17mgpu = Machine(GPU);
18
19m_2d = Machine(GPU);
20def same_point(Task task) {
21 return m_2d[*task.parent.processor(m_2d)];
22}
Figure A4: For the Circuit task, we show the mapper produced by the mapper agent at iteration 2.
1Task * GPU,OMP,CPU;
2Task calculate_new_currents GPU;
3Task update_voltages GPU;
4Region * * GPU FBMEM;
5Layout * * * C_order AOS Align==128;
6mgpu = Machine(GPU);
7
8m_2d = Machine(GPU);
9def same_point(Task task) {
10 return m_2d[*task.parent.processor(m_2d)];
11}
Figure A5: For the Circuit task, we show the mapper produced by the mapper agent at iteration 10.
1Task * GPU,OMP,CPU;
2Region * * GPU FBMEM;
3Region * * * SOCKMEM,SYSMEM;
4Layout * * * F_order SOA;
5mgpu = Machine(GPU);
6
7def block1d(Task task) {
8 ip = task.ipoint;
9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
10}
11
12IndexTaskMap task_2 block1d;
13
14m_2d = Machine(GPU);
15def same_point(Task task) {
16 return m_2d[*task.parent.processor(m_2d)];
17}
Figure A6: For Solomonik’s algorithm, we show the mapper produced by the mapper agent at iteration 2.
1Task * GPU,OMP,CPU;
2Region * * GPU FBMEM;
3Region * * * SOCKMEM,SYSMEM;
4Layout * * * C_order SOA No_Align;
5mgpu = Machine(GPU);
6
7def block1d(Task task) {
8 ip = task.ipoint;
9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
10}
11
12IndexTaskMap task_1 block1d;
13
14def cyclic1d(Task task) {
15 ip = task.ipoint;
16 linearize = ip[0] * 2 + ip[1];
17 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
18}
19
20IndexTaskMap task_1 cyclic1d;
21
22def cyclic2d(Task task) {
23 ip = task.ipoint;
24 linearize = ip[0] + ip[1] * 2;
25 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
26}
27
28IndexTaskMap task_1 cyclic2d;
29
30def linearize3D(Task task) {
31 ip = task.ipoint;
32 linearize = ip[0] + ip[1] + ip[2];
33 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
34}
35
36IndexTaskMap task_1 linearize3D;
37
38def linearize2D(Task task) {
39 ip = task.ipoint;
40 linearize = ip[0] * 2 + ip[2];
41 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
42}
43
44IndexTaskMap task_1 linearize2D;
45IndexTaskMap task_2 block1d;
46IndexTaskMap task_2 cyclic1d;
47IndexTaskMap task_2 cyclic2d;
48IndexTaskMap task_2 linearize3D;
49IndexTaskMap task_2 linearize2D;
50IndexTaskMap task_3 block1d;
51IndexTaskMap task_3 cyclic1d;
52IndexTaskMap task_3 cyclic2d;
53IndexTaskMap task_3 linearize3D;
54IndexTaskMap task_3 linearize2D;
55IndexTaskMap task_5 block1d;
56IndexTaskMap task_5 cyclic1d;
57IndexTaskMap task_5 cyclic2d;
58IndexTaskMap task_5 linearize3D;
59IndexTaskMap task_5 linearize2D;
60
61m_2d = Machine(GPU);
62def same_point(Task task) {
63 return m_2d[*task.parent.processor(m_2d)];
64}
Figure A7: For Solomonik’s algorithm, we show the mapper produced by the mapper agent at iteration 10.